Solve by the substitution method.
7x-9y=2
4x+32=y

Answer by amoresroy(333)

Given:
Equation 1:
7x-9y=2
Equation 2:
4x+32=y
Substitute the value of y in Equation 2 to Equation 1
Equation 1
7x-9y = 2
7x-9(4x+32) = 2
7x-36x-288 = 2
-29x = 2 + 288
-29x = 290
x = -10
Since y = 4x+32
y = 4(-10) + 32
y = -8